Calibri (Body) Wrap Text General a -- - 圈 $ v% , Merge & Center Conditional Format Cel Formatting as Table Styles high vol
1. Calculate the dead time for the proportional counter based on the two-source measurement. Explain why maintaining geometry is so important here. 2. Calculate the MDA (in epm) for alpha counting and for alpha/beta counting with the proportional counter.
Calibri (Body) "" Wrap Text General a -- - 圈 $ v% , Merge & Center Conditional Format Cel Formatting as Table Styles high voltag cpm Beta Sheet1+
Answer #1

The two source method requires you to take a total of three readings S1 (Source 1 alone) , S2(Source 2 alone) and S3 (combining the two sources)

\tau = \frac{X(\sqrt{1-Z})}{Y} \ \ \ \ \ \ \ \ \ \ (1)

where mim2 mbm12

Y = m_1m_2(m_{12}+m_b) - m_bm_{12}(m_1 + m_2)

Z =\frac{ Y(m_1+m_2-m_12-m_b)}{X^2}

  m1, m2 and m12 represent the corresponding observed rates. Also let nb and mb be the true and measured background rates

It is important that the geometries are kept the same as this method is a comparative method.

2)

The MDA is the minimum detectable activity of a counter.

MDA 271+4.65 t.Ef

where Rb is the counting rate , Ef is the efficiency of the detector and t is the count time